OpenStudy (mysticfury):

Can someone please explain to me how i can write a system of inequalities to represent a shaded part on my graph? I only have the graph to work with. Where should i start?

OpenStudy (mysticfury):

@3mar @AloneS @KamiBug @pooja195

OpenStudy (phi):

do you know what area has to be shaded ?

OpenStudy (mysticfury):

the graph that i have shows the 2 lines, were they intersect and the shaded region

OpenStudy (phi):

ok In that case, you need to figure out the equation of each line. Pick one of the lines. Can you find two points on the line (that are easy to read off the graph ) ?

OpenStudy (mysticfury):

yes i can. the first line has the points 3,1 and 0,4

OpenStudy (phi):

ok, find the slope (change in y divided by change in x) we already know 4 is the y-intercept (from the point 0,4 ) the equation will be y = mx + 4 where m is the slope you find.

OpenStudy (mysticfury):

okay so y = -1x + 4

OpenStudy (phi):

yes, or just y= -x+4 (we don't usually bother to write the 1) next, is the point (0,0) part of the shaded region ?

OpenStudy (mysticfury):

umm well the second line passes through the point 0,0

OpenStudy (phi):

Maybe I should have asked: is the shaded region "below" the line ?

OpenStudy (mysticfury):

yes the shaded region is below the first line

OpenStudy (phi):

also, is at the line solid or dotted ?

OpenStudy (mysticfury):

the first line is solid

OpenStudy (phi):

shaded below means we want all y values "beneath" the line y < -x+4 (if the line was dashed we use a less than) \( y \le -x+4 \) if the line is solid (we use less than or equal) as a check, point 0,0 is beneath the line, so the relation should be true: putting in 0 for both x and y we get 0 <= 0+4 0<= 4 which is true. So we have the first equation we now do the second line the same way: 1) find two points 2) find the slope 3) find the y-intercept 4) decide if we want the y's above (use > or >=) or below (use < or <=) the line

OpenStudy (mysticfury):

okay so for the second line, its dotted, the shaded is above the line, as for the points, the only ones that i can see are 3,1 . 0,0 . and -3, -1

OpenStudy (phi):

you only need two points to define a line. 0,0 and 3,1 look easy change in y is 1-0 change in x is 3-0 and the slope is ⅓ the y-intercept is 0 if we did not have a nice point like 0,0, we would find the y-intercept like this: write: y= ⅓ x + b using any point, say, 3,1 put in 3 for x and 1 for y 1= ⅓ * 3 + b 1= 1+b add -1 to both sides to get 1-1= 1-1+b or 0 = 0+b 0= b and b = 0 the equation is y= ⅓ x + 0 but of course we don't bother to write +0 y= ⅓ x we want the points above the line i.e. we want all the y's bigger than the line's y so change the = to >

OpenStudy (mysticfury):

so y < -x + 4 and y > 1/3x is the system of inequalities?

OpenStudy (phi):

almost. I thought the first line was solid. the "solid" means the line is part of the solution we use \( \le \) See attached for what this system looks like.

OpenStudy (mysticfury):

yes that looks just like the graph i have. and i know its supposed to have the line underneath but i wasnt sure how to make it on here.

OpenStudy (mysticfury):

Thank you so much for your help!!! :D

OpenStudy (phi):

btw, people use <= for \( \le \) you read it as "less than or equal"

OpenStudy (mysticfury):

ah okay that works

OpenStudy (3mar):

I am sorry @MysticFury for late! The electricity was off! With @phi, you are in good hands! Thanks for him for good explanation.